Solve the following system of equations by any method you wish
X + y = 5
X + 2z = 6
Y + 3z = 4

Answer by ankor@dixie-net.com(22740)

X + y = 5
X + 2z = 6
Y + 3z = 4
:
Rewrite it like this
x + y + 0z = 5
x + 0y + 2z = 6
0x + y + 3z = 4
:
Subtract the 3rd equation from the 1st equation
x + y + 0z = 5
0x + y + 3z = 4
--------------------subtraction eliminates y
x - 3z = 1
x = 3z + 1
:
Substitute (3z+1) for x in the 2nd equation
3z + 1 + 2z = 6
5z = 6 - 1
5z = 5
z = 1
:
Find x using x = 3z + 1
x = 3(1) + 1
x = 4
:
Find y using the 1st equation: x + y = 5
4 + y = 5
y = 5-4
y = 1
:
I'll let you check these values in the original equations